OpenStudy (anonymous):

sin^4 x + 2sin^2 x cos^2 x + cos^4 x = 1 prove the identity

myininaya (myininaya):

\[(\sin^2(x)+\cos^2(x))^2=(1)^2=1\]

OpenStudy (anonymous):

(sin^2 x + cos^2 x)^2 = 1

OpenStudy (anonymous):

what did you guys do?

myininaya (myininaya):

its a perfect square

myininaya (myininaya):

\[a^2+2ab+b^2=(a+b)^2\]

OpenStudy (anonymous):

how did I not see that?

myininaya (myininaya):

tricks and stuff

OpenStudy (anonymous):

thanks :)